. A spherical drop of oil has a mass of 20.5 ug and a radius of 1.74 mm. Given that the volume of a sphere is 4 3 pr 3 , calculate the density of the oil in kg/m3
Answer #1

r = 1.74 mm

= 1.74*10^-3 m

volume = (4/3)*pi*r^3

= (4/3)*3.14*(1.74*10^-3 m)^3

= 2.21*10^-8 m^3

mass = 20.5 ug

= 2.05*10^-5 g

Now use:

density = mass / volume

= (2.05*10^-5 g)/(2.21*10^-8 m^3)

= 928 g/m^3

= 0.928 Kg/m^3

Answer: 0.928 Kg/m^3